Workers: Miyo, Akutsu
Set calibration factors to the photosensor model based on the Simon's report on March 7 2018 (and posted on 15).
- PS1: Y: offset: -8402 (counts), gain: -1/173 (mm/counts)
- PS2: X: offset: -10317, gain: -1/234
- PS3: Z: offset: -8403, gain: -1/176
- Note that they have full meaning at the room temperature; will be changed in the cryogenic temperature.
- Did the final check before the full closing of the IYC; will be closed tomorrow, and then pumping down.
[Ongoing] Need cares of water springs from anchor bolts for oplev pylons.
[Ongoing] Need to confirm what are the scratch lines made by the measurement by the KEK team last week.
- Complete the calibration of the oplev; the arm length was reported by Simon, so....
- Check the balance of the suspended WAB with some wires and sensors are attached.
- Check how the damping would be now after the replacement of the damping magnets.
- Check the reproducibility of the alignment of the suspension among several modes.
- Check heights of several points.
- Take care of water spring??
- Two burndy connectors, vacuum compatible.
Stages outside the chamber to ease wiping our legs.
Good rail(?) to move the WAB suspension in the cryostat in the main beam direction.
Chain block? to support the load in the chamber.
- A clock to know the time easily.
- M6 screw set.